A car driver traveling at 12 m/sec. sees a stalled vehicle 17 meters ahead. He steps on the brake to decelerate the car. a) At what rate of his deceleration so that it stops just before hitting the stalled vehicle? b) At what time just before hitting the stalled vehicle.


1
Expert's answer
2021-02-05T05:03:15-0500

a)


"s=\frac{v^2-v_0^2}{2a}\to a=\frac{v^2-v_0^2}{2s}=\frac{0^2-12^2}{2\cdot17}=-4.24(m/s^2)" . Answer


b)


"v=v_0-at\to t=\frac{v_0-v}{a}=\frac{12-0}{4.24}=2.83(s)" . Answer